soybean meal is 12% protein:how many pounds of each should be mixed together in order to get 240 lbs mixture that is 9% protein.. how many lbs cornmeal should there be? how many lbs soybean should there be?


Answer by checkley77(12569)

.12x+.06(240-x)=.09*240
.12x+14.4-.06x=21.6
.06x=21.6-14.4
.06x=7.2
x=7.2/.06
x=120 pounds of soybeans is used.
240-120=120 pounds of cornmeal is used.
Proof:
.12*120+.06*120=.09*240
14.4+7.2=21.6
21.6=21.6